# Anticipating Object State Changes

### v1

@@ Abstract
Anticipating object state changes in images and videos is a challenging problem whose solution has important implications in vision-based scene understanding, automated monitoring systems, and action planning. In this work, we propose the first method for solving this problem. The proposed method predicts object state changes that will occur in the near future as a result of yet unseen human actions. To address this new problem, we propose a novel framework that integrates learnt visual features that represent the recent visual information, with natural language (NLP) features that represent past object state changes and actions. Leveraging the extensive and challenging Ego4D dataset which provides a large-scale collection of first-person perspective videos across numerous interaction scenarios, we introduce new curated annotation data for the object state change anticipation task (OSCA), noted as Ego4D-OSCA. An extensive experimental evaluation was conducted that demonstrates the efficacy of the proposed method in predicting object state changes in dynamic scenarios. The proposed work underscores the potential of integrating video and linguistic cues to enhance the predictive performance of video understanding systems. Moreover, it lays the groundwork for future research on the new task of object state change anticipation. The source code and the new annotation data (Ego4D-OSCA) will be made publicly available [^1] https://projects.ics.forth.gr/cvrl/osca/.
@@ 1 Introduction
When observing human-object interactions and object manipulations, we humans can effortlessly reason on and anticipate changes in object states  [6, 21, 3]. Imagine, for example, that while preparing the table for a dinner, somebody brings a bottle of wine. Even before opening it, we can fairly easily infer that in the near future, the bottle will be “opened”, and glasses will be “filled”. Recognizing and anticipating object states and their changes is crucial for any entity that interacts with objects, because the state of an object significantly affects its physical and functional properties and, therefore, plays a decisive role in further activity understanding, reasoning, and task planning.
@@
While it is almost effortless for humans, the capability of predicting object state changes still lies beyond the competencies of any current AI-powered system [38, 36]. Understanding object states and their changes in the context of actions relates to visual object perception, action/activity recognition, and object state estimation. While these individual capabilities have been heavily researched by the computer vision community, surprisingly, the problem of anticipating changes in object states remains unexplored. However, recognizing and anticipating object state changes would be an important competence of AI-powered agents that need to reason about human activities and/or perform task planning [33, 37].
Figure 1: We introduce a new problem, that of anticipating object state changes in videos of procedural activities. At a decision point which is the start time of the next, yet unobserved action, the goal is to predict whether there will be a state change related to the next active object and, if yes, estimate its type. An object state change is defined as the transition from a pre-state (initial) to a post-state (final) that occurs at the Point of No Return (PNR) time during a state-modifying action.
@@
Researchers have only recently started to focus on methods for the representation and understanding of object state changes in videos in the context of state-modifying actions [24, 28, 26], which can also be seen as transformations [31], while existing benchmarks largely ignore object state changes and focus on traditional types of annotations related to object type, location, or shape, attributes, affordances, and human actions. The problem of Object State Classification (OSC) is defined as the multi-class recognition of an object’s state in a still image [15, 13], or the initial and final object states in videos that demonstrate state-modifying actions [14]. The binary object state change classification variant is also related to the detection of whether or not a state change has occurred in a short video clip [7, 14].
@@
In this work, we take one step beyond Object State Classification (OSC) by introducing the new task of visual object state change anticipation (OSCA) in videos. OSCA focuses on the multi-class prediction of the state change to occur during the yet unseen, next action in an instructional video. Specifically, as shown in Fig. 1, at a certain decision point in time, noted as the start of the next, yet unobserved action, we aim to predict the type of state change of the next active object. The state change happens at the “Point of No Return” (PNR) [14]. Our approach may also predict that no state change will occur to the next active object in the forthcoming action which implies the anticipation of a no-state-modifying action. Our approach to tackle this new task relies on the cumulative history of recognized preceding actions and object states, up to the decision point where this prediction is performed. The integration of visual information concerning the recent past and textual description of past actions and states is a key idea to model this historical context effectively.
To assess the performance of our method for OSCA, we augment the annotations of the large-scale and challenging Ego4D dataset [14] to include labels for the initial and the final object states for state-modifying actions. This results in the Ego4D-OSCA, a variant of the Ego4D dataset that will become available to the vision community.
In summary, this work makes the following contributions:
@@
  • We introduce the new problem (OSCA) of anticipating the state change of the next active object in instructional videos.
  • We propose the first solution to OSCA, leveraging visual and language features for modeling the history of performed actions, object states and their changes.
@@
  • We introduce Ego4D-OSCA, a new benchmark for evaluating solutions to the OSCA problem.
@@ 2 Related Work
Object states capture the dynamic aspects of object appearance and/or functionality and are subject to visually perceivable changes, as a result/effect of state-modifying actions. They are also known as object fluents related to changeable object attributes [16, 2, 17]. Object State Change Anticipation (OSCA) is a new problem introduced in this paper. As such, to the best of our knowledge, there is no prior work that addresses this problem. Below, we review research on the closely-related problems of object state classification in images and the interplay of object state estimation and human action recognition in videos.
@@
Object State Classification in Images: Object states are typically considered as a special subset of “visual attributes”, i.e. visual concepts that are related to the physical and functional properties of objects [15]. Object states and their changes are perceivable by humans and should be perceivable by AI-enabled agents [8]. The majority of the approaches that address the attribute classification task follow a similar approach to that of object classification by training a convolutional neural network with discriminative classifiers on annotated image datasets [25]. Few works focus explicitly on state classification [13], while most of them rely on the same assumptions used for the attribute classification task. A prominent research direction to tackle this task refers to zero-shot learning that gained considerable attention in recent years due to its practical significance in real-world applications, mitigating the problem of collecting and learning training data for a very large number of object classes [34]. One prevalent zero-shot learning approach involves the utilization of semantic embeddings to represent objects and their attributes in a low-dimensional space [32]. The work in [12] leverages Knowledge Graphs (KGs) and semantic knowledge in the context of zero-shot object classification. In a similar vein, the work in [11] combines KGs and Large Language Models (LLMs) in order to address object-agnostic state classification. A recent work by [24] focuses on the recognition of object states based on the concept of compositional generation of novel object-state images, while the recently proposed method in [28] introduced a novel conditioned diffusion model that focuses on generating temporally consistent and physically plausible images of actions and object state transformations based on an input image and a text prompt describing the targeted transformation.
@@
Object State Change Estimation and Action Recognition in Videos: Object state changes have been considered a meaningful information source in the task of video-based human action understanding and recognition (HAR). Human actions frequently induce modifications in the states of associated objects, thereby influencing the overall dynamics of action sequences. Understanding the causal relationships between actions and object states has the potential to enrich our comprehension of action dynamics throughout long complex activities in procedural/instructional videos. Therefore, in HAR, object state changes are often considered complementary attributes to the visual representation of actions. These changes are typically derived within the visual domain through the utilization of explicit models for object detection and state estimation [9, 26, 18], or indirect modeling of object states based on general scene changes resulting from action execution [1, 5, 4]. Several methods exploit object states implicitly to estimate the type of action performed. The work in [1] was among the first to propose a method to automatically discover object states and the associated manipulation actions from videos by leveraging a discriminative clustering framework that jointly models the temporal order of object states and manipulation actions. The work in [17] explored the recognition of object fluents (changeable attributes of objects) and tasks (goal-oriented human activities) in egocentric videos using a hierarchical model that represents tasks as concurrent and sequential object fluents. Moreover, [19] focuses on understanding human actions within videos by analyzing complex interactions across multiple interrelated objects by recognizing different states changes of these objects. In [27, 26] a multi-task self-supervised framework is proposed that allows the temporal localization of object state changes and state-modifying actions in uncurated web videos. Furthermore, [35] introduced the novel VidOSC approach for understanding object state changes by segmenting object parts related to those changes in videos from an open-world perspective. A recently proposed framework [22] can recognize object-centric actions by relying only on the initial and final object states. The model is also able to generalize across unseen objects and different video datasets. The method proposed in [23] aims at disentangling visual embeddings that distinctly represent object states alongside identities, enabling effective recognition and generation of novel object-state compositions through a compositional learning framework. Finally, the InternVideo [7] video foundation model was adapted to tackle the tasks of object state change classification and action anticipation in the context of the Ego4D Challenges.
3 Methodology
Figure 2: Overview of the proposed baseline framework for the object state change anticipation task. The proposed framework anticipates object state changes by integrating real-time visual data and a historical record of past actions and object states changes.
@@ 3 Methodology
The proposed framework, depicted in Fig. 2, draws inspiration from the efficacy of amalgamating visual and lexical information for semantic action/activity encoding. Pursuant to this objective, it adopts a three-stream architecture. Within this design, a visual encoding module is tasked with capturing the visual attributes of the ongoing actions, while two lexical-based encoders are employed to extract the semantic nuances from a procedural-oriented representation of past actions and object states. The framework orchestrates the fusion of these distinct representations towards a unified objective: the anticipation of the next object state. This anticipatory task entails the estimation of the forthcoming state in which the object of interest will reside during the subsequent action. By integrating visual and lexical cues, the framework endeavors to holistically capture the underlying dynamics and contextual intricacies governing object-state transformations across sequential action.
@@
As illustrated in Fig. 2, the proposed framework consists of two primary components: (a) the current action and object state estimation module, and (b) the object state anticipation module, depicted within the thin-dotted rectangle. Our contribution resides in (a) the conceptualization of this framework and (b) the development of the object state anticipation module. Concerning the latter, it encompasses the following constituent components.
Visual Encoder: In the visual encoding module, we employ a lightweight visual encoder consisting of a single-branch bidirectional long short-term memory (BiLSTM) component followed by a multi-layer perceptron (MLP). We selected this simplified design for the visual encoder due to its primary objective of temporally encoding the enduring relationships among encoded short-term segments extracted from the input video. Our model relies on an external pretrained human action recognition model, such as SlowFast [10] or TSN [30] to provide encodings of short-term spatio-temporal dependencies between the frames inside a single segment.
@@
Learning Objective: The learning objective for training the model was exclusively focused on evaluating the anticipated state estimate. This deliberate choice stemmed from the desire to prioritize the accurate prediction of object states, which constituted the primary objective of the study. This objective was formulated using the cross-categorical entropy loss, which is well-suited for multi-class classification tasks, such as predicting object states across different categories:
@@
During training, the proposed framework leverages oracle action and state detectors to provide the action and state history, respectively, for each clip in the dataset (see Fig. 2). These detectors estimate the current action and object state observed in the clip, serving as ground truth annotations for training purposes. However, it is important to note that in a complete inference mode where the model is deployed for real-world applications, there is a requirement for current action and current object state recognition models to provide input to the framework. Consequently, our proposed model is solely tasked with the learning objective of next object anticipation during training, focusing exclusively on predicting the future state of the object. By decoupling the training and inference phases in this manner, the model is able to effectively learn the dynamics of object state transitions without the added complexity of simultaneously predicting the current action.
4 Experimental Setup
Figure 3: Six types of state-modifying actions related to two types of state transitions/changes (“deform” and “remove”), that are included in videos of the Ego4D dataset. Each action and state change occurrence is represented by a triplet of frames (pre-state, PNR, post-state). Each state change type is associated with various actions taking place in different environments and diverse scenarios, which emphasizes the complexity and the challenges introduced when dealing with the object state change anticipation (OSCA) problem.
The proposed Ego4D-OSCA dataset: We introduce Ego4D-OSCA as a new partition of the large-scale Ego4D dataset that aims to serve as a benchmark for the assessment of methods for object state change anticipation. The volume and diversity of the Ego4D dataset make Ego4D-OSCA a very challenging dataset for OSCA, as shown in Fig. 3. Ego4D-OSCA is tailored for the long-term activity (LTA) prediction benchmark, which aims to forecast the sequence of activities that will unfold in future video frames. Due to an ongoing challenge, the official test set for this benchmark has not yet been released, prompting us to re-purpose the validation split as a stand-in test set. To enrich the LTA benchmark, we have integrated object state annotations extracted from the dataset as follows.
The original Ego4D dataset does not include annotations for the specific state labels of individual video frames. Instead, annotations about state changes are provided, which relate to entire video segments. Additionally, the dataset includes annotations for bounding boxes and object classes across seven critical frames within each video segment. These frames are temporally centered around the occurrence of the state change that occurs within each video segment. Based on this information, we super-annotate certain critical frames of each video segment with state-related labels as follows. For each video segment, we annotate the initial and final frames as $pre\_X$ and $post\_X$, respectively, where $X$ denotes the label of the state change. Furthermore, in line with the semantic implications of these changes, we establish three pairs of state changes. Each pair is constructed under the premise that the first action is the inverse of the second concerning the resulting state change. For instance, if $X$ and $Y$ represent inverse state changes, then the labels $pre\_X$ and $post\_Y$ are considered samples of identical states. A similar correspondence applies between $pre\_Y$ and $post\_X$. For example, the states $pre\_remove$ and $post\_deposit$ are considered identical, since $remove$ and $deposit$ constitute a pair of inverse state changes. Table 1 delineates the specifics of these super-annotated states. Morever, Table 2 provides important dataset statistics. By incorporating these detailed annotations, Ego4D-OSCA offers researchers a comprehensive platform to explore and refine methods for anticipating object behavior and activity sequences in egocentric video contexts.
OSC
activate
deactivate
deposit
remove
construct
deconstruct
deform
other
Pre
pre
pre
pre
pre
pre
pre
pre
pre
activate
deactivate
deposit
remove
construct
deconstruct
deform
other
Post
post
post
post
post
post
post
post
post
activate
deactivate
deposit
remove
construct
deconstruct
deform
other
Table 1: The super-annotated state change labels and the corresponding pre-/post-state labels of a video segment, where the state modifying action occurs. The pairs activate-deactivate, deposit-remove, and construct-deconstruct constitute pairs of inverse state change actions. Frame state labels that correspond to the same state are depicted with the same colour.

### v2

@@ Abstract
In this work, we introduce (a) the new problem of anticipating object state changes in images and videos during procedural activities, (b) new curated annotation data for object state change classification based on the Ego4D dataset, and (c) the first method for addressing this challenging problem. Solutions to this new task have important implications in vision-based scene understanding, automated monitoring systems, and action planning. The proposed novel framework predicts object state changes that will occur in the near future due to yet unseen human actions by integrating learned visual features that represent recent visual information with natural language (NLP) features that represent past object state changes and actions. Leveraging the extensive and challenging Ego4D dataset which provides a large-scale collection of first-person perspective videos across numerous interaction scenarios, we introduce an extension noted Ego4D-OSCA that provides new curated annotation data for the object state change anticipation task (OSCA). An extensive experimental evaluation is presented demonstrating the proposed method’s efficacy in predicting object state changes in dynamic scenarios. The performance of the proposed approach also underscores the potential of integrating video and linguistic cues to enhance the predictive performance of video understanding systems and lays the groundwork for future research on the new task of object state change anticipation. The source code and the new annotation data (Ego4D-OSCA) will be made publicly available [^1] https://projects.ics.forth.gr/cvrl/osca/.
Figure 1: We introduce the new problem of anticipating object state changes in videos of procedural activities, noted OSCA. The decision point in the timeline is placed at the onset of the next anticipated, yet unobserved, human action. The objective is to accurately predict, at this point, the object state change class that will occur, if any, during the subsequent, yet unobserved action. This involves understanding the dynamics of past and current interactions and how they will affect the object’s state. An object state change (e.g. deform) refers to a functional and possibly to a physical change in an object’s attributes/properties. It is realized based on the transition from a pre-state (initial) to a post-state (final) occurring at the Point of No Return (PNR) time during an object state-modifying action (e.g. cut fish fillet using a knife).
@@ 1 Introduction
When observing human-object interactions, we can effortlessly reason about and anticipate changes in object states  [6, 32, 3]. Imagine, for example, that while preparing the table for a dinner, somebody brings a bottle of wine. Even before opening it, we can infer that in the near future, the bottle will be “opened”, and glasses will be “filled”. Recognizing and anticipating object states and their changes is crucial for any entity that interacts with objects because the state of an object significantly affects its physical and functional properties and plays a decisive role in activity understanding, reasoning, and task planning.
@@
While it is almost effortless for humans, the capability of predicting object state changes still lies beyond the competencies of current AI-powered systems [62, 58]. Understanding object states and their changes in the context of interactions relates to several challenging tasks, such as visual object perception, next-active object prediction, action recognition and anticipation, and object state estimation, that have been well-explored by the research community. Surprisingly, the problem of anticipating object state changes remains undefined and unexplored. However, recognizing and anticipating object state changes would be an important ability of AI-powered agents toward understanding human activities and task planning [53, 61].
@@
The problem of Object State Classification (OSC) is defined as the multi-class recognition of an object’s state in a still image [22, 17], or the initial and final object states in videos that demonstrate state-modifying actions [18]. The binary object state change classification variant is also related to the detection of state change occurrences in a short video clip [7, 18]. Researchers have only recently started to focus on methods for the representation and understanding of object state changes in videos in the context of state-modifying actions [38, 44, 41], which can also be seen as transformations [50]. Existing benchmarks largely ignore object state changes and focus on traditional types of annotations related to object type, location, or shape, attributes, affordances, and human actions.
@@
In this work, we take one step beyond the Object State Classification and the Action Anticipation tasks by introducing the new task of Object State Change Anticipation (OSCA) in videos. OSCA focuses on the multi-class prediction of the state change that will occur on an object during the yet unseen, next action in an instructional video. Specifically, as shown in Fig. 1, at a certain decision point in time that is at the start of the next, yet unobserved action, we aim to predict the object state change class that will occur. The object state change will occur at the “Point of No Return” (PNR) timestamp during the next action [18]. To tackle this new task, we introduce a new formulation leveraging on the cumulative history of the textual description of recognized preceding actions and object states, up to the decision point where this prediction is performed. The integration of the former information with visual information concerning the recent past is a key idea to model this historical context effectively. Our approach may also predict that no object state change will occur in the forthcoming action which implies the anticipation of a no-state-modifying action. We evaluate the proposed approach toward the newly proposed OSCA problem to establish baseline results for automated change state anticipation in long instructional videos and to also investigate the potential impact and effectiveness of the extracted information in other related vision-based anticipation tasks. We assess the performance of our proposed method for the OSCA problem and show initial results based on a proposed extension of the popular Ego4D video dataset. We hence build on the large-scale and challenging Ego4D dataset [18] which provides ego-centric videos by augmenting the available annotation data with labels for the object state changes based on the initial and the final object states for any state-modifying actions in a subset of videos related to the Hand & Object Interactions benchmark  [^2] https://ego4d-data.org/docs/benchmarks/hands-and-objects/. This results in the Ego4D-OSCA, a variant of the Ego4D dataset that will become available to the computer vision community.
Thus, the main contributions of this work are:
@@
  • The introduction of the new problem of anticipating an object state change (OSCA) that will occur in the next, yet unseen, action in instructional videos.
@@
  • The introduction of the Ego4D-OSCA dataset, a new benchmark for evaluating solutions to the OSCA problem based on a subset of the Ego4D video dataset.
  • The proposal of the first approach to tackle OSCA which integrates visual and language features to model the history of performed actions, object states, and their changes. We also present initial baseline results.
Figure 2: Examples of modifying actions from the “deform” and “remove” object state change classes. Each example is represented by a triplet of frames (pre-state, PNR, post-state). A state change is associated with various actions taking place in diverse environments and scenarios, emphasizing the complexity and challenges introduced in the OSCA problem.
@@ 2 Related Work
Object states capture dynamic aspects of object appearance and/or functionality and are subject to visually perceivable changes, as a result/effect of state-modifying actions. They are also known as object fluents related to changeable object attributes [24, 2, 25]. Since there is no prior work on the problem of Object State Change Anticipation (OSCA) that we introduce, we review the literature on the closely related topics of action anticipation, next-active object anticipation, and object state classification in images as well as the interplay between object state estimation and action recognition in videos.
2.1 Object State Classification/Recognition
@@
Object State Classification in Images: Object states are typically considered as a special subset of “visual attributes”, i.e. visual concepts that are related to the physical and functional properties of objects [22]. Object states and their changes are perceivable by humans and should be perceivable by AI-enabled agents [10]. The majority of the attribute classification approaches follow a similar approach to that of object classification by training a convolutional neural network with discriminative classifiers on annotated image datasets [40]. Few works focus explicitly on state classification [17], while most of them rely on the same assumptions used for the attribute classification task. A prominent direction to tackle this task refers to zero-shot learning. It gained considerable attention in recent years due to its practical significance in real-world applications, mitigating the problem of collecting and learning training data for a very large number of object classes [54]. One such prevalent approach involves the utilization of semantic embeddings to represent objects and their attributes in a low-dimensional space [51]. The work in [16] leverages Knowledge Graphs (KGs) and semantic knowledge in the context of zero-shot object classification. In a similar vein, the work in [15] combines KGs and Large Language Models (LLMs) to address object-agnostic state classification. A recent work by [38] focuses on object state recognition based on the compositional generation of novel object-state images, while the method in [44] introduced a novel conditioned diffusion model that focuses on generating temporally consistent and physically plausible images of actions and object state transformations based on an input image and a text prompt describing the targeted transformation.
@@
Object State Change Estimation & Action Recognition in Videos: Object state changes have been considered a meaningful information source in video-based human action understanding and recognition (HAR). In HAR, object state changes are often considered complementary attributes to the visual representation of actions. These changes are typically derived within the visual domain through the utilization of explicit models for object detection and state estimation [11, 41, 26], or indirect modeling of object states based on general scene changes resulting from action execution [1, 5, 4]. Several methods exploit object states implicitly to estimate the type of action performed. The work in [1] was among the first to propose a method to automatically discover object states and the associated manipulation actions from videos by leveraging a discriminative clustering framework that jointly models the temporal order of object states and manipulation actions. The work in [25] explored the recognition of object fluents (changeable attributes of objects) and tasks (goal-oriented human activities) in egocentric videos using a hierarchical model that represents tasks as concurrent and sequential object fluents. Moreover, [27] focuses on understanding human actions within videos by analyzing complex interactions across multiple interrelated objects by recognizing different state changes of these objects. In [43, 41] a multi-task self-supervised framework is proposed that allows the temporal localization of object state changes and state-modifying actions in uncurated web videos.
Furthermore, [56] introduced the novel VidOSC approach for understanding object state changes by segmenting object parts related to those changes in videos from an open-world perspective. A recently proposed framework [36] can recognize object-centric actions by relying only on the initial and final object states. The model can also generalize across unseen objects and different video datasets. The method proposed in [37] aims at disentangling visual embeddings that distinctly represent object states alongside identities, enabling effective recognition and generation of novel object-state compositions through a compositional learning framework. Finally, the InternVideo [7] video foundation model was adapted to tackle the tasks of object state change classification and action anticipation in the context of the Ego4D Challenges.
2.2 Action & Next-Active Object Anticipation
Action anticipation involves predicting the label of an action that is expected to occur in the future but has not yet been started/observed [20, 60, 52]. This challenge has been studied in both egocentric [18, 8] and exocentric [39, 21] videos, with the latter becoming increasingly popular in recent years. Short-term action anticipation [19, 35, 31] focuses on predicting actions or events in the immediate future, whereas long-term action anticipation [59, 30] extends to predicting actions or events over a longer period, ranging from several seconds to minutes.
In the context of human-object interactions in videos, active objects [33] and next-active objects [14] refer to specific items that are involved in ongoing or anticipated actions. The active object is the item that a person is currently interacting with within the video. In contrast, the next-active object is the item predicted to be used in the near future, based on the current interaction [23, 29, 9]. Although not yet in use, it is likely to become involved in subsequent actions. These concepts are crucial in video analysis for understanding and predicting human behavior, as they help anticipate the sequence of actions and interactions within a scene. The concept of next-active object anticipation has also been the subject of the short-term anticipation challenge in [18] and is described as the next object that will be touched by the user (either with their hands or with a tool) to initiate an interaction. Several methods have been proposed in this challenge for the solution of this problem [45, 46, 34, 31].
Figure 3: Two of the nine object state change classes that have been super-annotated in the proposed Ego4D-OSCA dataset, ”deposit” and ”remove,” are illustrated. Pre-/post-state labels for two state-modifying actions are shown as distinct video segments. ”Deposit” and ”remove” are inverse state changes, where the pre-deposit matches the post-remove, and the pre-remove matches the post-deposit, as indicated by frames and shapes of the same color.
Anticipating the state change of an object involves predicting how the object’s condition or form will alter as it becomes involved in an activity, whether it is currently active or will become active in the near future. This process goes beyond merely identifying which object will be used next(next-active); it focuses on understanding how the object’s state will evolve during or after its involvement in the interaction. This anticipatory process requires analyzing the current interaction and understanding the transformations that occur as objects are used. By predicting these state changes, we gain insights into how objects will behave as they become active or next-active, enhancing the ability to interpret the sequence of events and interactions in a video.
Figure 4: Annotation pipeline: Occlusions are checked in the pre- & post-frames. A threshold value for the BBOX area of $N$ square pixels (N=100) for each object annotation. Ego4D-SCOD benchmark data are used to automatically annotate the change states per clip.
3 Ego4D-OSCA Dataset
We introduce Ego4D-OSCA as a new partition of the large-scale Ego4D dataset that aims to serve as a benchmark for the assessment of methods for object state change anticipation. The volume and diversity of the Ego4D dataset make Ego4D-OSCA a very challenging dataset for OSCA, as shown in Fig. 2. Ego4D-OSCA is tailored from the long-term activity (LTA) prediction benchmark, which aims to forecast the sequence of activities that will unfold in future video frames. Due to an ongoing challenge, the official test set for this benchmark has not yet been released, prompting us to re-purpose the validation split as a stand-in test set.
Dataset annotation: To enrich the LTA benchmark, we have integrated object state annotations extracted from the dataset as follows. The original Ego4D dataset does not include annotations for the specific state labels of individual video frames. Instead, annotations about state changes are provided, which relate to entire video segments. Additionally, the dataset includes annotations for bounding boxes and object classes across seven critical frames within each video segment. These frames are temporally centered around the occurrence of the state change that occurs within each video segment. Based on this information, we super-annotate certain critical frames of each video segment with state-related labels as follows. For each video segment, we annotate the initial and final frames as $pre\_X$ and $post\_X$, respectively, where $X$ denotes the label of the state change. Furthermore, in line with the semantic implications of these changes, we establish three pairs of state changes. Each pair is constructed under the premise that the first action is the inverse of the second concerning the resulting state change. For instance, if $X$ and $Y$ represent inverse state changes, then the labels $pre\_X$ and $post\_Y$ are considered samples of identical states. A similar correspondence applies between $pre\_Y$ and $post\_X$. For example, the states $pre\_remove$ and $post\_deposit$ are considered identical, since $remove$ and $deposit$ constitute a pair of inverse state changes. Figure 3 delineates the specifics for these two super-annotated state change classes. The same condition is true for the state change classes of activate-deactivate and construct-deconstruct. The full set of pre- and post- object state pairs that constitute the target set of object state changes appear in the supplementary material. It is important to note that there exists a subset of video segments containing actions that do not induce state changes and, therefore, these segments are not considered for annotation.
By incorporating these detailed annotations, Ego4D-OSCA offers researchers a comprehensive platform to explore and refine methods for anticipating object behavior and activity sequences in egocentric video contexts. The Ego4D dataset offers eight distinct state change labels: activate, deactivate, deposit, remove, construct, deconstruct, deform, and other. However, we contend that there are actions that do not alter the state of an object. To address this, we propose adding a state change category called “No Object-State-Change (No OSC)”. This new class will help capture instances where actions occur without affecting the state of an object, thereby providing a more comprehensive framework for understanding and categorizing interactions.
Datasets Modalities
OSC related task
Actions per video
Samples
Obj. State Classes
Actions
Objects
Fire et al. [13] Videos Detection Single 490 17 - 13
ChangeIt [42] Videos Temporal Localization Single 34.428 - 44 -
HowToChange [55] Video & Text Temporal Localization Single 498.475 20 - 134
VSCOS [57] Video Segmentation Single 1.905 4 271 124
VOST [47] Video Segmentation Single 713 - - 155
Ego4D [18] Video & Text Detection & Classification Single 92.864 8 - 478
Ego4D-OSCA (Ours) Video & Text Anticipation Multiple 1610 9 1500 477
Table 1: Comparison with other image and video datasets that contain annotations related to object state changes. Modalities refer to the available data source for the object state change-related tasks.
Details on the annotation process: The annotation process for the state transitions is applied to the $pre\_Y$ and $post\_X$ frames in each video segment. Overall, the annotation process consists of the next 4 steps (a schematic representation of the annotation pipeline is shown in Figure 4). First, the PNR moment of the video segment being examined for annotation is compared to the PNR moment of the segment that has been previously annotated. If the PNR of the previously annotated segment is located after the PNR of the segment under examination, then the segment under examination is rejected. The reasoning behind this decision has to do with the learning of the segment features related to state transitions. This alignment of the two PNRs signifies that there is an overlap between the state transition actions of the two segments and therefore the feature learning becomes more challenging. Subsequently, it is examined if the object undergoing state change is occluded. If this is the case, the frame is rejected. Then, the bounding box area of the object is evaluated, and if it is below 100 square pixels—a threshold empirically chosen and commonly used in annotation tools like Voxel-51—the frame is discarded. Finally, the frame that has passed all the previous checks is annotated with the appropriate state label that pertains to the state transition action.
Dataset Statistics: The proposed Ego4D-OSCA dataset is compiled using a subset of the popular, large-scale Ego4D v2 dataset [18] that contains egocentric videos for a large variety of human daily living or work activities. In Table 1 we compare the proposed Ego4D-OSCA dataset with existing image and video datasets that also provide annotations related to object states. The target tasks performed using each of the datasets are also noted. Moreover, Figure 5 provides statistics about the state change classes. More dataset statistics can be found in the supplementary material.
Figure 5: Statistics for the Ego4D-OSCA dataset by object state change class, with 61.858 training and 31.846 testing clips.
4 Object State Change Anticipation - Baseline
@@
The proposed framework, depicted in Fig. 6, draws inspiration from the efficacy of combining visual and lexical information for semantic action/activity encoding. To achieve this, it adopts a three-stream architecture. Within this design, a visual encoding module is tasked with capturing the visual attributes of ongoing actions, while two lexical-based encoders are employed to extract the semantic nuances from a procedural-oriented representation of past actions and object states. The framework fuses these distinct representations towards the unified objective of anticipating the next object state. This task entails the estimation of the forthcoming state in which the object of interest will reside during the subsequent action. By integrating visual and lexical cues, the framework tries to holistically capture the underlying dynamics and contextual intricacies governing object-state transformations across sequential actions.
@@
As illustrated in Fig. 6, the proposed framework consists of two primary components: (a) the current action and object state estimation module, and (b) the object state anticipation module, depicted within the thin-dotted rectangle. Our contribution resides in (a) the conceptualization of this framework and (b) the development of the object state anticipation module. Concerning the latter, it encompasses some constituent components.
Figure 6: Overview of the proposed baseline framework for the object state change anticipation task. The proposed framework anticipates object state changes by integrating real-time visual data and a historical record of past actions and object state changes.
Visual Encoder: For this module, we employ a lightweight visual encoder consisting of a single-branch bidirectional long short-term memory (BiLSTM) component followed by a multi-layer perceptron (MLP). We selected this simplified design for the visual encoder based on the objective of temporally encoding the enduring relationships among encoded short-term segments extracted from the input video. Our model relies on an external pretrained human action recognition model, such as SlowFast [12] or TSN [49] to provide encodings of short-term spatio-temporal dependencies between the frames inside a single segment.
@@
Learning Objective: The objective for training the model was exclusively focused on evaluating the anticipated state estimate. This deliberate choice stemmed from the aim to prioritize the accurate prediction of object states, which was the study’s primary objective. This objective was formulated using the cross-categorical entropy loss, which is well-suited for multi-class classification tasks, such as predicting object states across different categories:
@@
During training, the proposed framework leverages oracle action and state detectors to provide the action and state history, respectively, for each clip in the dataset (see Fig. 6). These detectors estimate the current action and object state observed in the clip, serving as ground truth annotations for training purposes. However, it is important to note that for inference toward real-world applications, there is a requirement for current action and object state recognition models to provide input to the framework. Consequently, our proposed model is solely tasked with the learning objective of next-object anticipation during training, focusing exclusively on predicting the future state of the object. By decoupling the training and inference phases in this manner, the model can effectively learn the dynamics of object-state transitions without the added complexity of simultaneously predicting the current action.
5 Implementation, Experiments and Results